# SpecGen: Automated Generation of Formal Program Specifications via Large Language Models

### v2

@@ Abstract
To tackle this challenge, we introduce SpecGen, a novel technique for formal program specification generation based on Large Language Models (LLMs). Our key insight is to overcome the limitations of existing methods by leveraging the code comprehension capability of LLMs. The process of SpecGen consists of two phases. The first phase employs a conversational approach that guides the LLM to generate appropriate specifications for a given program, aiming to utilize the ability of LLM to generate high-quality specifications. The second phase, designed for where the LLM fails to generate correct specifications, applies four mutation operators to the model-generated specifications and selects verifiable specifications from the mutated ones through a novel heuristic selection strategy by assigning different weights of variants in an efficient manner. We evaluate SpecGen on two datasets, including the SV-COMP Java category benchmark and a manually constructed dataset containing 120 programs. Experimental results demonstrate that SpecGen succeeds in generating verifiable specifications for 279 out of 385 programs, outperforming the existing purely LLM-based approaches and conventional specification generation tools like Houdini and Daikon. Further investigations on the quality of generated specifications indicate that SpecGen can comprehensively articulate the behaviors of the input program.
@@ I Introduction
Yet, a practical challenge is the absence of documented formal specifications in most real-world software projects, since manually writing high-quality specifications is highly nontrivial. To alleviate the burden on software developers, several tools have been introduced for generating program specifications automatically [7, 8, 9], including Houdini [7] and Daikon [9], two most representative ones for Java programs. However, these tools rely heavily on predefined templates or grammars during the specification generation process. This often results in overly simplistic specifications that struggle to accurately capture the complex behaviors and functionalities of real-world programs. Consequently, the specifications generated by such tools usually lack practical significance in the actual software development process.
@@
To evaluate SpecGen, we conduct experiments on two datasets. We first evaluate SpecGen on the benchmark for the Java category of SV-COMP [16]. To further evaluate the performance of SpecGen on different kinds of programs, we constructed another dataset containing 120 Java programs with manually written ground-truth specifications. The selected programs are highly representative, encompassing different control-flow structures and various data structures to avoid any bias in our evaluation. We compared the performance of SpecGen on the dataset against three baselines: Houdini, Daikon, and Purely LLM-based methods. The results of our evaluation demonstrate that SpecGen significantly outperforms the baseline methods. SpecGen successfully generated verifiable specifications for 279 out of the total 385 programs, outweighing 218 for conversational specification generation, the best-performing Purely LLM-based method, and 98 for Houdini, the best-performing non-LLM method. An ablation study on mutations was also conducted, proving the effectiveness of all four types of mutation operators. Additionally, the results of evaluations on the heuristic selection strategy suggested that our strategy effectively improves the efficiency of SpecGen compared to the random selection strategy. Furthermore, a user study was conducted to evaluate the semantic quality of the generated specifications, illustrating the ability of SpecGen to accurately and comprehensively characterize program behaviors. The main contributions are summarized as follows:
@@
  • A dataset namely SpecGenBench with hand-written specifications by experts, facilitating follow-up research. Other than the established benchmark SV-COMP, we collected programs on a more diverse spectrum for deeper insights.
@@ II-B Motivation
Recently, large language models (i.e., LLMs) [23, 24] have exhibited powerful capacities in coding [15, 14]. The emergence of these models may greatly compensate for the limitations of traditional software analysis tools in code understanding. A significant amount of work attempts to leverage large language models in software engineering [25, 26, 27, 28, 29] and we have witnessed substantial progress brought about by the introduction of LLMs. Inspired by these works, in this work, we aim to leverage the large language models in the automated generation of formal program specifications to address the limitations of conventional templated-based approaches. From this perspective, we innovate our approach SpecGen, which generated three parts of specifications for the example presented in the right part of Figure 1. The first part is to describe method TwoSum() as a whole, specifying its pre-conditions and post-conditions. The specifications in lines a and b claim that the input array must not be null before and after the method is executed. The post-condition in line c specifies that the target value equals the sum of the two elements corresponding to the indexes stored in the returned array. The post-condition at line d specifies that there does not exist such a pair of elements that satisfies the constraint when the length of the returned array is zero. The functionality of method TwoSum() can be fully articulated by these generated specifications. Furthermore, the loop invariants in the second and third parts specify corresponding constraints that must be met within a certain range in the array. These generated specifications by SpecGen comprehensively describe the semantics of this function and their correctness is verified by the verifier.
@@ III-C1 Template Specification Mutation
As shown in Table I, we define four kinds of mutation operators including predicative, logical, comparative, and arithmetic. Each type of mutation corresponds to one type of operator supported by JML. LLMs perform well in capturing the overall semantics of a program, but they often make mistakes in grasping the fine-grained relationships between variables, resulting in incorrect operators used to describe the relationships between variables, which is why our mutation design is centered around the operators. When applying the mutation operator to a specification, it substitutes the operators of the corresponding type in the specification with another operator of the same type. For example, after applying a predicative mutation, a $\exists$ predicate within a specification may be substituted with $\forall$. Note that the mutation for a certain type of operator does not necessarily create only one mutated candidate. For example, the expression a <= b may be mutated to a < b or a - 1 <= b. If multiple types of mutations can be applied to a specification at the same time, we try to exhaust each combination of different types of mutations to get all potential variants. For instance, the expression x < n + 1 can be mutated to x <= n + 1 from the comparative type, x < n - 1 from the arithmetic type, or x <= n - 1 by combining them.
@@ IV-B Dataset
To comprehensively evaluate the effectiveness of SpecGen, following previous work [36], we first use an established dataset namely SV-COMP [16] benchmark for evaluation. Specifically, we used 265 class definitions in the Java category of SV-COMP benchmark and conducted the necessary modifications on part of these programs (referred to as Dataset SV-COMP hereinafter) for ease of evaluation. The remaining data in the benchmark cannot be applied for specification generation even with our modification. In particular, the verification with expected verification result FALSE (i.e., a counterexample to safety) has to be modified so that the program can exit properly, and those library calls specific to the competition settings (e.g. Verifier.nondetInt()) in the definitions are replaced with equivalent Java library calls so that they can be successfully compiled. However, after a deep analysis of the characteristics of the data from SV-COMP, we find that 88.7% programs are loop-free, indicating that the samples with more complex program structures cannot be covered by this dataset, inducing limitations on the evaluation. Also, very few datasets have been established specifically for specification generation tasks so far.
@@
To remedy this gap, we further collect another dataset containing 120 samples as a supplement where 20 programs (including the corresponding specifications) from the dataset constructed by Nilizadeh et al. [37] and 100 programs from LeetCode [38]. This dataset is namely SpecGenBench and we have made it public [17]. The selected programs are assured of the feasibility of expressing their behaviors as JML-specified verifiable specifications. These programs involve a variety of control flow structures and encompass multiple data structures such as arrays, strings, and other data structures supported by the Java library. They also cover a diverse set of specifications including post-conditions and loop invariants, involving both linear and nonlinear relationships between variables, making them representative of a broad spectrum of scenarios. They can be categorized into five categories according to their types of control flow structures [39]. Specifically, Sequential denotes the programs without branches or loops. Branched represents the loop-free programs that will contain branches like if-else or switch structures. Single-path Loop contains the simplest type of loop, with only one layer of loop structure without branches in their loop bodies. In contrast, Multi-path Loop denotes the loops that have branches in the loop bodies. Lastly, Nested Loop denotes the programs with multiple layers of loop structure where each layer may have a branch. The quantity for Sequential, Branched, Single-path Loop, Multi-path Loop and Nested Loop is 26, 23, 24, 26, and 21, respectively.
@@
To obtain the oracle specifications for the 100 Java programs from LeetCode, we follow a similar procedure in Nilizadeh et al. [37] with the help of human experts. Three experts with rich experience in formal verification were employed, to manually write specifications for each program. Each expert is required to write the specifications that can be successfully verified to describe the functionality and behavior of the program as accurately and comprehensively as possible. For a single program, if multiple experts have written verifiable specifications, another expert is responsible for selecting one of them as the oracle.
@@
As Daikon [9] requires a set of test suites instrumented into the source code to execute the code, we also manually write these test suites for it. Lastly, we instrument the dummy function calls at the top of each loop body in a program to ensure Houdini and Daikon can generate the specifications at these program points.
@@ IV-C Baselines
We select two conventional approaches and a purely LLM-based approach as the baselines for comparison.
@@
Apart from the above approaches used for specification generation, we further add the purely LLM-based approaches.
Purely LLM-based Approaches. They refer to those relying entirely on the large language model i.e., gpt-3.5 in our work to generate specifications (Section III-B). We compare SpecGen with few-shot setting and conversational setting. We set 0-shot, 2-shots, and 4-shots for few-shot comparison. The conversational setting is also based on 4-shot examples and the error message is provided for conversation. The other settings are the same with SpecGen.
@@ V-A RQ1: Comparison with Baselines
The experimental results are presented in Table II where Num. denotes the number of successfully handled programs, and Prob. denotes the average success probability. For the purely LLM-based approaches and SpecGen, the number of trials for each program is 10 times.
@@
Performance on SV-COMP. From Table II, we can find that on the SV-COMP dataset (265 programs in total), Houdini handled 56 programs, which is more than Daikon. However, both of them underperform LLM-based approaches even in LLM’s simplest setting i.e. 0-shot setting, which handled 81 programs, demonstrating the feasibility of employing LLMs for formal program specification generation. Among the LLM-based approaches with few-shot settings, as the number of given few-shot examples increases, the number of programs that LLM can generate verifiable specifications is also increased, substantiating the effectiveness of the few-shot examples. Based on the 4-shots examples in the initial prompt, the multi-turn conversational manner (in Section III-B) can further improve the performance, with 146 programs handled, compared to 94 programs of 4-shot LLM. Lastly, SpecGen outperforms the conversational-based approach, with the number of programs that generated verifiable specifications increasing from 146 to 179.
@@
Performance on SpecGenBench. For deeper insights into the performance of different approaches on different types of programs, we further investigate the results on SpecGenBench. Similar to SV-COMP, Houdini outperforms Daikon, especially in generating specifications for complex program structures such as loops, Houdini exhibits certain abilities. In terms of the LLM-based approaches, we find that both SpecGen and the conversational-based approach are competent in generating specifications for relatively simple programs such as sequential and branched. However, when it comes to generating more complex program structure specifications such as loop, SpecGen has a clear advantage. The benefits are from our designed mutation-based specification generation (Section III-C), which can correct the erroneous output of the large language model to generate the verifiable specifications. Although SpecGen can generate more accurate specifications for different loop structures, it has a relatively poor performance in generating verifiable specifications that the program has the nested loop.
@@
For a program, we further use the metric namely success probability to evaluate the probability of successful generation for this program by LLMs in 10 times trials due to the randomness inherent in the generated content by LLMs. We can observe that SpecGen achieves an overall success probability of 59.97%, which is significantly higher than the value of conversational manner i.e., 35.95%. It indicates that compared with the purely conversational manner, SpecGen is more reliable with higher probabilities and lower chances of randomness to generate verifiable specifications. From the results from SV-COMP and SpecGenBench, we can find that our proposed approach is orthogonal to different datasets.
@@
To the end, the Venn diagram of the programs in SpecGenBench for which SpecGen and baseline approaches successfully generated verifiable specifications is presented in Fig. 5. We merge the programs successfully handled by different purely LLM-based approaches together as Purely LLM-based in Fig. 5. It is noteworthy that SpecGen generates verifiable specifications for 14 programs that other baselines fail to generate, where 6 are from the Nested Loop, 3 are from the Multi-path Loop, and 5 are from the Single-path Loop. Further investigation of these 14 programs, we find that they are relatively complicated, making it hard for the baseline approaches to correctly generate. The generated specifications of different approaches for these programs are presented on our website [17]. We also present a case study to illustrate the successful/unsuccessful generation on the website.
RQ1: SpecGen significantly outperforms the baselines on two different datasets, generating verifiable specifications for 279 out of 385 programs, whereas the conversation-driven approach, Houdini and Daikon succeeds in 218, 98, and 72 respectively.

### v3

@@ Abstract
To tackle this challenge, we introduce SpecGen, a novel technique for formal program specification generation based on Large Language Models (LLMs). Our key insight is to overcome the limitations of existing methods by leveraging the code comprehension capability of LLMs. The process of SpecGen consists of two phases. The first phase employs a conversational approach that guides the LLM in generating appropriate specifications for a given program, aiming to utilize the ability of LLM to generate high-quality specifications. The second phase, designed for where the LLM fails to generate correct specifications, applies four mutation operators to the model-generated specifications and selects verifiable specifications from the mutated ones through a novel heuristic selection strategy by assigning different weights of variants in an efficient manner. We evaluate SpecGen on two datasets, including the SV-COMP Java category benchmark and a manually constructed dataset containing 120 programs. Experimental results demonstrate that SpecGen succeeds in generating verifiable specifications for 279 out of 385 programs, outperforming the existing LLM-based approaches and conventional specification generation tools like Houdini and Daikon. Further investigations on the quality of generated specifications indicate that SpecGen can comprehensively articulate the behaviors of the input program.
@@ I Introduction
Yet, a practical challenge is the absence of documented formal specifications in most real-world software projects, since manually writing high-quality specifications is highly nontrivial. To alleviate the burden on software developers, several tools have been introduced for generating program specifications automatically [7, 8, 9], including Houdini [7] and Daikon [9], two most representative ones for Java programs. However, these tools rely heavily on predefined templates or grammars during the specification generation process. As claimed by Molina et al. [10], the fixed templates involved result in a limited range of specifications covered, usually yielding overly simplistic specifications that struggle to capture the complex behaviors and functionalities of real-world programs accurately. This phenomenon poses non-negligible limitations for these tools, consequently hindering their applications in the actual software development process [11, 12, 13].
@@
To evaluate SpecGen, we conduct experiments on two datasets. We first evaluate SpecGen on the benchmark for the Java category of SV-COMP [20]. To further evaluate the performance of SpecGen on different kinds of programs, we constructed another dataset containing 120 Java programs with manually written ground-truth specifications. The selected programs are highly representative, encompassing different control-flow structures and various data structures to avoid any bias in our evaluation. We compared the performance of SpecGen on the dataset against multiple baselines. The results of our evaluation demonstrate that SpecGen significantly outperforms the baseline methods. SpecGen successfully generated verifiable specifications for 279 out of the total 385 programs, outweighing 247 for AutoSpec [21], the best-performing LLM-based approach, and 98 for Houdini, the best-performing non-LLM method. An ablation study on mutations was also conducted, proving the effectiveness of all four types of mutation operators. Additionally, the results of evaluations on the heuristic selection strategy suggested that our strategy effectively improves the efficiency of SpecGen compared to the random selection strategy. Furthermore, a user study was conducted to evaluate the semantic quality of the generated specifications, illustrating the ability of SpecGen to accurately and comprehensively characterize program behaviors. The main contributions are summarized as follows:
@@
  • A dataset named SpecGenBench, with hand-written specifications by experts, facilitating follow-up research. Other than the established benchmark SV-COMP, we collected programs on a more diverse spectrum for deeper insights.
@@ II-B Motivation
Recently, large language models (i.e., LLMs) [28, 29] have exhibited powerful capacities in coding [19, 18]. The emergence of these models may greatly compensate for the limitations of traditional software analysis tools in code understanding. A significant amount of work attempts to leverage large language models in software engineering [30, 31, 32, 33, 34] and we have witnessed substantial progress brought about by the introduction of LLMs. Inspired by these works, in this work, we aim to leverage the large language models in the automated generation of formal program specifications to address the limitations of conventional template-based approaches. From this perspective, we innovate our approach SpecGen, which generated three parts of specifications for the example presented in the right part of Figure 1. The first part is to describe the method TwoSum() as a whole, specifying its pre-conditions and post-conditions. The specifications in lines a and b claim that the input array must not be null before and after the method is executed. The post-condition in line c specifies that the target value equals the sum of the two elements corresponding to the indexes stored in the returned array. The post-condition at line d specifies that there does not exist such a pair of elements that satisfies the constraint when the length of the returned array is zero. The functionality of method TwoSum() can be fully articulated by these generated specifications. Furthermore, the loop invariants in the second and third parts specify corresponding constraints that must be met within a certain range in the array. These generated specifications by SpecGen comprehensively describe the semantics of this function and their correctness is verified by the verifier.
@@ III-C1 Template Specification Mutation
As shown in Table I, we define four kinds of mutation operators including predicative, logical, comparative, and arithmetic. Each type of mutation corresponds to one type of operator supported by JML. LLMs perform well in capturing the overall semantics of a program, but they often make mistakes in grasping the fine-grained relationships between variables, resulting in incorrect operators used to describe the relationships between variables, which is why our mutation design is centered around the operators. When applying the mutation operator to a specification, it substitutes the operators of the corresponding type in the specification with another operator of the same type. For example, after applying a predicative mutation, a $\exists$ predicate within a specification may be substituted with $\forall$. Note that the mutation for a certain type of operator does not necessarily create only one mutated candidate. For example, the expression a <= b may be mutated to a < b or a - 1 <= b. If multiple mutations can be applied to a specification at the same time, we try to exhaust each combination of different types of mutations to get all potential variants. Since the set of all potential variants of a certain template is determined, the exhaustive searching process is deterministic. For instance, the expression x < n + 1 can be mutated to x <= n + 1 from the comparative type, x < n - 1 from the arithmetic type, or x <= n - 1 by combining them.
@@ IV-B Dataset
To comprehensively evaluate the effectiveness of SpecGen, following previous work [41], we first use an established dataset, the benchmark of SV-COMP [20], for evaluation. Specifically, we used 265 class definitions in the Java category of SV-COMP benchmark and conducted the necessary modifications on part of these programs (referred to as Dataset SV-COMP hereinafter) for ease of evaluation. The remaining data in the benchmark cannot be applied for specification generation even with our modification. We made minimal modifications to SV-COMP programs to ensure they can be executed outside the SV-COMP environment. Specifically, the programs destined to trigger false assertions have to be modified so that the programs can exit properly. Also, those library calls specific to the competition settings (e.g. Verifier.nondetInt()) are replaced with equivalent Java library calls so that they can be successfully compiled. It is ensured that the semantics of the modified programs remain unchanged. As calculated by tool JaCoCo [42], these programs have an average line of code (LoC) of 22.51, along with an average cyclomatic complexity (CC) of 6.18. However, after a deep analysis of the characteristics of the data from SV-COMP, we find that 88.7% programs are loop-free, indicating that the samples with more complex program structures cannot be covered by this dataset, inducing limitations on the evaluation. Also, very few datasets have been established specifically for specification generation tasks so far.
@@
To remedy this gap, we further collect another dataset, SpecGenBench, containing 120 samples as a supplement where 20 programs (including the corresponding specifications) from the dataset constructed by Nilizadeh et al. [43] and 100 programs from LeetCode [44]. This dataset is publicly accessible [22]. The selected programs are assured of the feasibility of expressing their behaviors as JML-specified verifiable specifications. These programs involve a variety of control flow structures and encompass multiple data structures such as arrays, strings, and other data structures supported by the Java library. They also cover a diverse set of specifications including post-conditions and loop invariants, involving both linear and nonlinear relationships between variables, making them representative of a broad spectrum of scenarios. They can be categorized into five categories according to their types of control flow structures [45]. Specifically, Sequential denotes the programs without branches or loops. Branched represents the loop-free programs that will contain branches like if-else or switch structures. Single-path Loop contains the simplest type of loop, with only one layer of loop structure without branches in their loop bodies. In contrast, Multi-path Loop denotes the loops that have branches in the loop bodies. Lastly, Nested Loop denotes the programs with multiple layers of loop structure where each layer may have a branch. The quantity of programs for Sequential, Branched, Single-path Loop, Multi-path Loop and Nested Loop is 26, 23, 24, 26, and 21, respectively. Programs in SpecGenBench have an average LoC of 20.77 and an average CC of 6.60.
@@
To obtain the ground truth specifications for the 100 Java programs from LeetCode, we follow a similar procedure in Nilizadeh et al. [43] with the help of human experts. Three experts with rich experience in formal verification were employed, to manually write specifications for each program. Each expert is required to write the specifications that can be successfully verified to describe the functionality and behavior of the program as accurately and comprehensively as possible. For a single program, if multiple experts have written verifiable specifications, another expert is responsible for selecting one of them as the ground truth.
@@
As Daikon [9] requires a set of test suites instrumented into the source code to execute the code, we manually write these test suites for it. Specifically, we initialize a small test suite for a program. Then, Daikon is invoked based on the execution of this test suite to generate specifications. If the verifier cannot verify the generated specifications, we adopt a counterexample-guided strategy to include the counterexample produced by the verifier in the test suite. The procedure is repeated continuously until no new specifications are generated. As calculated by JaCoCo [42], the test suites achieve an average instruction coverage of 90.16%, branch coverage of 87.98%, and line coverage of 91.36%. Lastly, we instrument the dummy function calls at the top of each loop body in a program to ensure Houdini and Daikon can generate the specifications at these program points.
@@ IV-C Baselines
We select two conventional approaches and several LLM-based approaches as the baselines for comparison.
@@
Apart from the conventional approaches for specification generation, we further add the LLM-based approaches.
Few-shot LLM. They refer to the large language model i.e., gpt-3.5 with few-shot settings in our work to generate specifications. Under the few-shot settings, the LLM is queried only once to obtain the final result. We set 0-shot, 2-shot, and 4-shot for few-shot comparison.
Conversational. It refers to the generation technique described in Section III-B. Conversational generation iteratively queries the LLM to refine its results, with error information provided to the LLM as feedback on each iteration. The conversational setting is based on 4-shot examples. Other settings remain the same with SpecGen.
AutoSpec [21]. It is a recent technique for specification generation combining LLMs and static analysis. The approach comprises three steps: code decomposition, hierarchical specification generation, and specification validation. AutoSpec first decomposes the input program into its components, upon which a hierarchy graph is built. For each component, AutoSpec queries the LLM for corresponding specifications respectively. Eventually, specifications for all components are combined to obtain the overall result, which is presented to a verifier for correctness verification. The progress is repeated iteratively until the result is successfully verified.
@@ V-A RQ1: Comparison with Baselines
The experimental results are presented in Table II where Num. denotes the number of successfully handled programs, and Prob. denotes the average success probability. For the LLM-based approaches, the number of trials for each program is 10 times.
@@
Performance on SV-COMP. From Table II, we can find that on the SV-COMP dataset (265 programs in total), Houdini handled 56 programs, which is more than Daikon. However, both underperform LLM-based approaches even in LLM’s simplest setting i.e. 0-shot setting, which handled 81 programs, demonstrating the feasibility of employing LLMs for formal program specification generation. Among the LLM-based approaches with few-shot settings, as the number of given few-shot examples increases, the number of programs that LLM can generate verifiable specifications is also increased, substantiating the effectiveness of the few-shot examples. Based on the 4-shots examples in the initial prompt, the multi-turn conversational manner (in Section III-B) can further improve the performance, with 146 programs handled, compared to 94 programs of 4-shot LLM. Combining LLM-generated results and static analysis techniques, AutoSpec achieves enhanced results with 10 more programs handled. Lastly, SpecGen outperforms all baseline approaches, with the number of programs that generated verifiable specifications increasing to 179.
@@
Performance on SpecGenBench. For deeper insights into the performance of different approaches on different types of programs, we further investigate the results on SpecGenBench. Similar to SV-COMP, Houdini outperforms Daikon, especially in generating specifications for complex program structures such as loops, Houdini exhibits certain abilities. In terms of the LLM-based approaches, we find that SpecGen, AutoSpec, and the conversational approach are all competent in generating specifications for relatively simple programs such as sequential and branched. AutoSpec also demonstrates impressive ability on programs with Single-path Loops, benefiting from the code decomposition technique adopted. However, when it comes to generating specifications for programs with more complex structures such as Multi-path Loops and Nested Loops, SpecGen has a clear advantage. The benefits are from our designed mutation-based specification generation (Section III-C), which can correct the erroneous output of the large language model to generate the verifiable specifications. Although SpecGen can generate more accurate specifications for different loop structures, it has a relatively poor performance in generating verifiable specifications for programs with nested loops.
@@
For LLM-based approaches, we further use success probability to evaluate the probability of successful generation for a program in 10 times trials due to the randomness inherent in the generated content by LLMs. We can observe that SpecGen achieves an overall success probability of 59.97%, which is significantly higher than the values of conversational generation and AutoSpec (35.95% and 46.13% respectively). AutoSpec is not equipped with any feedback-refine mechanism. Although conversational generation can refine the intermediate results through conversation with LLM, there still exists detailed errors that cannot be fixed, since root causes of errors are difficult for LLMs to reason. Compared to these approaches, SpecGen, equipped with conversational generation and mutation-based generation, is more reliable with higher probabilities and lower chances of randomness to generate verifiable specifications. From the results from SV-COMP and SpecGenBench, we can find that our proposed approach is orthogonal to different datasets.
@@
Presented in Fig. 5 is the Venn diagram of the programs in SpecGenBench for which SpecGen and baseline approaches successfully generated verifiable specifications. It is noteworthy that SpecGen generates verifiable specifications for 7 programs that other baselines fail to yield, where 5 are from the Nested Loop, with the rest 2 from the Single-path Loop. Further investigation of these programs reveals that they are relatively complicated, making it hard for the baseline approaches to be correctly generated. The generated specifications of all approaches for these programs can be accessed on our website [22]. where a case study is also presented to illustrate the successful/unsuccessful generation.
Overall, it takes on average 7.07 verifier calls for the conversational approach per execution, whereas the figure is 16.19 and 15.51 for AutoSpec and SpecGen, respectively. Under our experimental settings, the maximum number of rounds for conversation is set to 10, so conversational generation naturally finishes execution faster than the other two approaches, but with relatively poor performance. In comparison, AutoSpec and SpecGen take more verifier calls to filter the LLM-generated results and produce more reliable results. Specifically, SpecGen displays a slight advantage over AutoSpec, indicating that SpecGen can achieve better performance compared to AutoSpec in the same or even shorter period of time.
RQ1: SpecGen outperforms all the baselines on two different datasets, generating verifiable specifications for 279 out of 385 programs, with the highest success probability among all LLM-based approaches. In comparison, the number of programs handled by Daikon, Houdini, Conversational generation, and AutoSpec is 72, 98, 218, and 247, respectively.